Find three consecutive even integers such that the sum of the smallest and the largest is 36.
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

Find three consecutive even integers such that the sum of the smallest and the largest is 36.
---
1st: 2x-2
2nd: 2x
3rd: 2x+2
------
Equation:
2x-2 + 2x+2 = 36
4x = 36
x = 9
---
1st: 2x-2 = 18-2 = 16
2nd: 18
3rd: 20
============
Cheers,
Stan H.
================

Answer by lwsshak3(11628)

Find three consecutive even integers such that the sum of the smallest and the largest is 36.
**
let x=smallest even integer
x+2=second consecutive even integer
x+4=third and largest consecutive integer
..
x+(x+4)=36
2x+4=36
2x=36-4=32
x=16
x+2=18
x+4=20
ans:
Three consecutive integers are: 16, 18 and 20